M is the midpoint of line segment AB. The coordinates of M are (1, 3) and the coordinates of A are (5, -3). Find the coordinates of B.
Answer by stanbon(75887)

M is the midpoint of line segment AB. The coordinates of M are (1, 3) and the coordinates of A are (5, -3). Find the coordinates of B.
----------------
Let the coordinates of B be (x,y)
-----
Equations:
(x+5)/2 = 1
x+5 = 2
x = -3
---------------
(y-3)/2 = 3
y-3 = 6
y = 9
----------------
Ans: B is (-3,9)
-------------------
Cheers,
Stan H.
---------------